3. (21 pts) Compute the area of the region enclosed by the curves y = -2° + 5 and y = - 9x + 20.
Answer #1

fl= y = -x² +52 and g(0) = y=x2-gar 420 (upper function - action)-llower loan functionare ty=-22h (upper function) 51 2 y= re- then Soy bonbs s = +8+ PŁ- = tot sit ose- - Osz = 3 DT82- 917 99)- SEIt öst- = (ob-ort ;-)-6001 – sti + -) = (ove-ac29e+ęce